# Revisiting Financial Sentiment Analysis: A Language Model Approach

### v1

@@ Abstract
Financial Sentiment Analysis (FSA) traditionally relies on human-annotated sentiment labels to infer investor sentiment and forecast market movements. However, inferring the potential market impact of words based on human-perceived intentions is inherently challenging. We hypothesize that the historical market reactions to words, offer a more reliable indicator of their potential impact on markets than subjective sentiment interpretations by human annotators. To test this hypothesis, a market-derived labeling approach is proposed to assign tweet labels based on ensuing short-term price trends, enabling the language model to capture the relationship between textual signals and market dynamics directly. A domain-specific language model was fine-tuned on these labels, achieving up to an 11% improvement in short-term trend prediction accuracy over traditional sentiment-based benchmarks. Moreover, by incorporating market and temporal context through prompt-tuning, the proposed context-aware language model demonstrated an accuracy of 89.6% on a curated dataset of 227 impactful Bitcoin-related news events with significant market impacts. Aggregating daily tweet predictions into trading signals, our method outperformed traditional fusion models (which combine sentiment-based and price-based predictions). It challenged the assumption that sentiment-based signals are inferior to price-based predictions in forecasting market movements. Backtesting these signals across three distinct market regimes yielded robust Sharpe ratios of up to 5.07 in trending markets and 3.73 in neutral markets. Our findings demonstrate that language models can serve as effective short-term market predictors. This paradigm shift underscores the untapped capabilities of language models in financial decision-making and opens new avenues for market prediction applications.
@@ 1 Introduction
  • Market-Driven Labeling: We replace subjective sentiment labels with market-derived labels that directly reflect ensuing price movements, enabling the model to learn from actual market behavior rather than inferred sentiment.
  • Context-Enhanced Prompt Tuning: By embedding key market information into textual prompts, our approach expands the feature space of language models, providing the model with the context required for precise market predictions.
  • Comprehensive Strategy Evaluation: Model predictions are translated into trading signals and rigorously assessed through backtesting over three years of Bitcoin market data, covering bullish, bearish, and neutral market regimes.
@@ 3 Related Works
Sentiment analysis in financial markets emerged from the growing recognition that market psychology plays a crucial role in market movements [9]. As Natural Language Processing (NLP) techniques advanced, they enabled the extraction of sentiments and emotions from large volumes of unstructured text, providing a means to quantify market psychology [10, 11]. Social media platforms, particularly Twitter, became valuable sources of real-time insights into investor sentiment, which could be used to predict financial market behavior. Sentiment analysis, or opinion mining, had already been established as a method within NLP for extracting sentiments, making them well-suited for quantifying investor psychology and applying these signals to financial markets. The power of sentiment analysis in forecasting market trends, volatility, volume, and risk contributed to the birth of FSA [3, 2]. The primary goal of FSA has been to utilize NLP methods to capture investor sentiment and psychology, providing valuable insights for predicting market movements [2]. Over time, the field has evolved, and various methods have been developed to refine sentiment classification, improve prediction accuracy, and integrate more nuanced factors such as market context. In the subsequent sections, we will explore the current state of the art in FSA, examine existing gaps, and highlight how this study aims to address these challenges.
@@ 3.1 Financial Sentiment Analysis Scope
FSA spans a wide range of methods, including lexicon-based approaches [12], word representations [13], and pre-trained language models [14, 15, 16, 17]. It also draws on various data sources such as news headlines [18], social media posts [19], and financial documents [3], with applications in volatility prediction, price prediction, and trend forecasting [2]. The core goal of FSA is to quantify investor sentiment from textual data to inform decision-making in financial markets. In this study, we focus on predicting Bitcoin’s short-term market movements by quantifying sentiment from Bitcoin-related tweets. This task has been addressed by various studies, but we consider PreBit [20] to be the state-of-the-art in this area. Unlike many other studies, PreBit advances the field by passing the actual tweet representations, rather than their sentiment class labels, to the trend prediction model. This allows the model to learn patterns between the words in the tweets and the market movements they ought to predict. By retaining the tweet’s representation, the model can identify nuanced relationships between specific words and their effect on market behavior, bridging the gap between textual sentiment and market outcomes.
@@ 4.4 Signal Generation and Backtesting Strategies
Table 1: Evaluation of base models Base-1, and Base-2, (respectively FinBERT and CryptoBERT) and proposed methods on the 2020 5-fold cross-validation dataset ($E_{a}$) and event-sampled Bitcoin data ($E_{b}$). The results for $E_{a}$ are the averages of metrics across 5 folds, with standard deviations representing fold-to-fold variation. For the base models, $E_{a}$ results reflect variations across the different data folds.

### v2

# Market-Derived Financial Sentiment Analysis: Context-Aware Language Models for Crypto Forecasting

@@ Abstract
Financial Sentiment Analysis (FSA) traditionally relies on human-annotated sentiment labels to infer investor sentiment and forecast market movements. However, inferring the potential market impact of words based on their human-perceived intentions is inherently challenging. We hypothesize that the historical market reactions to words, offer a more reliable indicator of their potential impact on markets than subjective sentiment interpretations by human annotators. To test this hypothesis, a market-derived labeling approach is proposed to assign tweet labels based on ensuing short-term price trends, enabling the language model to capture the relationship between textual signals and market dynamics directly. A domain-specific language model was fine-tuned on these labels, achieving up to an 11% improvement in short-term trend prediction accuracy over traditional sentiment-based benchmarks. Moreover, by incorporating market and temporal context through prompt-tuning, the proposed context-aware language model demonstrated an accuracy of 89.6% on a curated dataset of 227 impactful Bitcoin-related news events with significant market impacts. Aggregating daily tweet predictions into trading signals, our method outperformed traditional fusion models (which combine sentiment-based and price-based predictions). It challenged the assumption that sentiment-based signals are inferior to price-based predictions in forecasting market movements. Backtesting these signals across three distinct market regimes yielded robust Sharpe ratios of up to 5.07 in trending markets and 3.73 in neutral markets. Our findings demonstrate that language models can serve as effective short-term market predictors. This paradigm shift underscores the untapped capabilities of language models in financial decision-making and opens new avenues for market prediction applications.
@@ 1 Introduction
  • Market-derived Labeling: We replace subjective sentiment labels with market-derived labels that directly reflect ensuing price movements, enabling the model to learn from actual market behavior rather than inferred sentiment.
  • Contextualized Prompt-tuning: By embedding key market information into textual prompts, our approach expands the feature space of language models, providing the model with the context required for precise market predictions.
  • Trading Strategy Evaluation: Model predictions are translated into trading signals and rigorously assessed through backtesting over three years of Bitcoin market data, covering bullish, bearish, and neutral market regimes.
@@ 3 Related Works
Sentiment analysis in financial markets emerged from the growing recognition that market psychology plays a crucial role in market movements [9]. As Natural Language Processing (NLP) techniques advanced, they enabled the extraction of sentiments and emotions from large volumes of unstructured text, providing a means to quantify market psychology [10]. Social media platforms, particularly Twitter, became valuable sources of real-time insights into investor sentiment, which could be used to predict financial market behavior. Sentiment analysis, or opinion mining, had already been established as a method within NLP for extracting emotional responses [11], making them well-suited for quantifying investor psychology and applying these signals to financial markets. The power of sentiment analysis in forecasting market trends, volatility, volume, and risk contributed to the birth of FSA [3, 2]. The primary goal of FSA has been to utilize NLP methods to capture investor sentiment and psychology, providing valuable insights for predicting market movements [2]. Over time, the field has evolved, and various methods have been developed to refine sentiment classification, improve prediction accuracy, and integrate more nuanced factors such as market context. In the subsequent sections, we will explore the current state of the art in FSA, examine existing gaps, and highlight how this study aims to address these challenges.
@@ 3.1 Financial Sentiment Analysis Scope
FSA spans a wide range of methods, including lexicon-based approaches [12], word representations [13], and pre-trained language models [14, 15, 16, 17, 18]. It also draws on various data sources such as news headlines [13], social media posts [19], and financial documents [3], with applications in volatility prediction, price prediction, and trend forecasting [2]. The core goal of FSA is to quantify investor sentiment from textual data to inform decision-making in financial markets. In this study, we focus on predicting Bitcoin’s short-term market movements by quantifying sentiment from Bitcoin-related tweets. This task has been addressed by various studies, but we consider PreBit [20] to be the state-of-the-art in this area. Unlike many other studies, PreBit advances the field by passing the actual tweet representations, rather than their sentiment class labels, to the trend prediction model. This allows the model to learn patterns between the words in the tweets and the market movements they ought to predict. By retaining the tweet’s representation, the model can identify nuanced relationships between specific words and their effect on market behavior, bridging the gap between textual sentiment and market outcomes.
@@ 4.4 Signal Generation and Backtesting Strategies
Table 1: Evaluation of base models Base-1, and Base-2, (respectively FinBERT [15] and CryptoBERT [16]) and proposed methods on the 2020 5-fold cross-validation dataset ($E_{a}$) and event-sampled Bitcoin data ($E_{b}$). The results for $E_{a}$ are the averages of metrics across 5 folds, with standard deviations representing fold-to-fold variation. For the base models, $E_{a}$ results reflect variations across the different data folds.